# Efficient Model Editing with Task Vector Bases: A Theoretical Framework and Scalable Approach

### v1

@@ 6.1 Models and Datasets
This paper includes both computer vision and natural language processing experiments. In the vision experiments, we use CLIP (Radford et al., 2021) and ViT (Dosovitskiy, 2020) models from OpenCLIP (Ilharco et al., 2021), along with the following 8 datasets: MNIST (LeCun et al., 1998), EuroSAT (Helber et al., 2019), Cars (Krause et al., 2013), GTSRB (Stallkamp et al., 2012), RESISC45 (Cheng et al., 2017), DTD (Cimpoi et al., 2014), SUN397 (Xiao et al., 2010), and SVHN (Netzer et al., 2011). The language experiments are based on RoBERTa (Liu, 2019) models, tested on a medium scale benchmark with 12 tasks (Panigrahi et al., 2023; He et al., 2024) and a larger scale collection of 70 tasks, the latter with further details provided in Appendix E.
@@ 6.2 Empirical Verification of Theoretical Results
We primarily focus on Section 4.2 and examine the relationship between the loss gap and key constants.
@@
To verify how task vector similarity $\epsilon$ impacts the performance, we conduct the experiment shown in Figure 2. We merge the MNIST task vector with each of the other task vectors, all having similar norms ranging from $[2,3)$ (see Table 7 in the Appendix for details), and set the scaling coefficient $\alpha$ to $0.5$. In this setting, we approximately control all constants in Section 4.2, including $L$, $\alpha$, and $C$, and observe that highly similar tasks, such as digit classification in MNIST, SVHN, and GTSRB, lead to larger loss gaps or worse performance for MNIST compared to less related tasks.
@@
We provide additional evidence that both $C$ and $\epsilon$ must be constrained for the success of task vector addition. In Table 2, we collected task vectors for 8 tasks from two CLIP checkpoints pretrained with different hyperparameters. From this table, we observe that successful task addition reveals the identity of the task vectors. For optimal merging performance, we should only add task vectors fine-tuned from the same checkpoint, as any mixture of task vectors from different checkpoints will cause a significant performance drop. The above empirical observation consolidates our 4.1 that task vectors should reside in the same fine-tuning regime. To elaborate, from Figure 3, although all $\epsilon$ values in the pink box are very low, task addition still fails due to the large $C$ value. From Figure 4, we see that with different hyperparameters, the two pretrained models are situated in two local convex basins, and the distance between the two checkpoints is much larger than the task vectors (for more discussion, see Section D.1). Thus, if we create task vectors by subtracting the wrong pretrained checkpoint, the large $C$ value leads to the failure of task addition.
@@ 6.3 Vision Experiments with Task Vector Bases
For a detailed analysis of the behavior of task vector bases methods, we first fix a small subset of tasks where the separation between digit classification (MNIST, SVHN) and satellite image classification (EuroSAT, RESISC45) is obvious and matches the optimal clustering result that reduces the memory storage by half. With only $4$ task vectors, we can enumerate all possible partitions of tasks and understand the position of the merging performance using this natural partition.
@@
After saving task vector bases from MNIST, SVHN, EuroSAT, and RESISC45 in Section 6.3, we test generalization metrics on unseen tasks in our vision benchmark. We set a similarity threshold to only use the most similar task vector bases for GTSRB and DTD generalization in Table 5, with $\texttt{merge}_{\text{out}}$ set to Task Arithmetic. Since we use only one basis for generalization, this reduces to tuning $\alpha$ for the selected basis on the validation set. For GTSRB, which involves traffic sign recognition, we expect overlap with the digit classification skills in MNIST and SVHN. In Table 5, we observe that the performance of the MS-ER column, corresponding to the bases method, is generally higher than the M-S-E-R column, where full task arithmetic flexibility is retained. For OOD generalization, using all task-specific information related to the target task proves beneficial. However, since DTD is not related to any stored bases, its OOD generalization performance is almost equivalent to the pretrained model. This validates Section 4.2 and shows that when an OOD task is not too far away from seen tasks, leveraging bases knowledge is more effective than relying solely on a task vector trained on a single task, with the added benefit of memory savings.
@@ 6.4.1 Results for Large Scale Benchmark
We collect 70 natural language understanding tasks from Huggingface (Wolf, 2019) datasets that can be reformulated as classification problems. With a larger number of tasks, we can conduct a stress test to determine how much we can compress the memory budget using Algorithm 1. Following Gao et al. (2020); He et al. (2024), we design appropriate prompt templates and labels for masked language modeling fine-tuning of RoBERTa-base, ensuring that each fine-tuned model’s performance exceeds the majority vote baseline. As a result, each of the collected models is meaningful, and since all tasks can be evaluated with accuracy, the merged metrics are directly comparable.
We first examine the result of clustering in Figure 6, where we show the task cosine similarity matrix ordered by the spectral clustering results. We plot two matrices: $k=3$, which represents the number of clusters selected based on the largest eigenvalue gap, and $k=10$, which further partitions tasks into more fine-grained groups. Consequently, The properties of the task vectors change according to the prompt template difference. When $k=3$, the tasks are split based on the number of input sentences, and when $k=10$, there are more fine-grained clusters that reflect the task distribution, such as the separation of bAbI-NLI tasks, which reformulate tasks from Weston et al. (2015) as NLI problems, and other types of NLI tasks. In summary, the generated task clusters in step 2 of Figure 1 are reasonable, as they reflect certain interpretable skills learned by all models. And when we use $k=3$ as the algorithm suggested, we reduce the storage of 70 task vectors into 3 bases, saving 95.71% of the memory budget.
Table 6: Merged accuracy/F1 and storage cost on the $12$-task language benchmark. First two rows contain numbers in He et al. (2024).
@@ 6.4.2 Results for Medium Scale Benchmark
(L&S) He et al. (2024) divides model merging into two steps: first, localizing the task information by learning a sparse vector mask, and second, stitching sparse models with mask normalization, ensuring that $\alpha$ is a probability vector like 4.5. Compared to other model merging methods, He et al. (2024) offers the dual advantage with significantly improved multitask performance due to reduced conflict from localization, and drastically lowered memory costs by saving only a small proportion of nonzero weights. Thus, to maximally boost all metrics, we investigate how to integrate L&S into our bases creation framework which requires two steps of merging. As described in Section 5.1, after clustering sparse localized task vectors, we can directly apply L&S as $\texttt{merge}_{\text{in}}$ within each cluster. For $\texttt{merge}_{\text{out}}$, it requires special design and we propose two modifications to L&S.
@@
The second approach, Stitch Twice, first localizes tasks and stitches by mask normalization within cluster. Afterward, in $\texttt{merge}_{\text{out}}$, we round within-cluster-normalized masks as updated binary task masks, and use them as input for the second iteration of stitching across all $T$ tasks. Formally, in the first iteration, for $l\in[d]$, the $l$-th entry of the mask for task $i$ is $\gamma_{i}^{l}\leftarrow\texttt{round}(\gamma_{i}^{l}/\sum_{j=1}^{N}\gamma_{j}^{l})$, and the second iteration of mask update is $\gamma_{i}^{l}\leftarrow\gamma_{i}^{l}/\sum_{j=1}^{T}\gamma_{j}^{l}$. Inspired by the improvements shown in Figure 7, this two-step process automatically readjusts task vector weights with appropriate task relationships learned from clustering. Unlike Mask Sharing, Stitch Twice requires $T$ copies of sparse models, since each task retains its own mask. However, in Table 6, Stitch Twice improves generalization performance from L&S (and more competitive baselines, see Section E.1) while also surprisingly reducing memory costs due to increased mask sparsity from double normalization.
@@ 7 Conclusion and Future Work
In this work, we revisited task vectors as a practical approach for model editing. By providing a theoretical foundation for task arithmetic, we addressed key gaps in understanding its underlying mechanisms and identified the conditions for its success. Building on these insights, we introduced a scalable framework that reduces the memory footprint of task vector arithmetic through clustered task bases, significantly improving efficiency while maintaining high performance, and demonstrated the flexibility of bases arithmetic. Looking forward, we foresee several future directions. First, the creation of our task vector bases depends on the task relationship, and there are additional opportunities to improve memory efficiency from hardware perspectives. Additionally, by clustering, we assume that there are disjoint groups of tasks for the initial task vectors. Therefore, it might be interesting to further decompose and reorganize task vectors to boost performance for more complicated task relationships.
@@ Acknowledgment
SZ, YH, and HZ are partially supported by an NSF IIS grant No. 2416897. HZ would like to thank the support from a Google Research Scholar Award. WY is supported by a gift from AWS AI for research in Trustworthy AI. MY is partly supported by MEXT KAKENHI Grant Number 24K03004 and by JST ASPIRE JPMJAP2302. We are thankful to Wei Hu, Yuki Takezawa, Jose Renato Restom for helpful discussions. The views and conclusions expressed in this paper are solely those of the authors and do not necessarily reflect the official policies or positions of the supporting companies and government agencies.
@@
which means that $\theta^{*}$ is within the fine-tuning regime and satisfies the local smoothness assumption. Hence, if $x\sim\mathcal{D}_{i}$,
${\mathcal{L}}_{i}(\theta^{*})-\mathcal{L}_{i}(\theta_{i})$ $\theta^{*}-\theta_{i},\frac{\partial\mathcal{L}_{i}(x,\theta_{i})}{\partial{\theta}}\right>+\frac{L_{i}}{2}\left\lVert\theta^{*}-\theta_{i}\right\rVert^{2}$
@@
[…]
$\leq\sum\limits_{j=1}^{T}\alpha_{j}^{2}\left\lVert\tau_{j}\right\rVert^{2}+2\sum_{j\neq k}\alpha_{j}\alpha_{k}\left|\left<\tau_{j},\tau_{k}\right>\right|+2\sum\limits_{j=1}^{T}\alpha_{j}\left|\left<\tau_{j},\tau_{i}\right>\right|+\left\lVert\tau_{i}\right\rVert^{2}$
$=\sum\limits_{j=1}^{T}\alpha_{j}^{2}\left\lVert\tau_{j}\right\rVert^{2}+2\sum_{j\neq k}\alpha_{j}\alpha_{k}\left|\left<\tau_{j},\tau_{k}\right>\right|+2\sum\limits_{j\neq i}^{T}\alpha_{j}\left|\left<\tau_{j},\tau_{i}\right>\right|+2\alpha_{i}\left\lVert\tau_{i}\right\rVert^{2}+\left\lVert\tau_{i}\right\rVert^{2}$
$\leq\left(\sum\limits_{j=1}^{T}\alpha_{j}^{2}+2\alpha_{i}+1\right)C+2\left(\sum_{j\neq k}\alpha_{j}\alpha_{k}+\sum\limits_{j\neq i}^{T}\alpha_{j}\right)\epsilon C.$
To further upper bound the coefficients of the above two terms, we have
$\sum\limits_{j=1}^{T}\alpha_{j}^{2}+2\alpha_{i}+1\leq\sum_{j=1}^{T}\alpha_{j}+2+1=4,$
and
$42\left(\sum_{j\neq k}\alpha_{j}\alpha_{k}+\sum\limits_{j\neq i}^{T}\alpha_{j}\right)\leq 2\left(\left(\sum_{j=1}^{T}\alpha_{j}\right)\left(\sum_{k=1}^{T}\alpha_{k}\right)+\sum\limits_{j=1}^{T}\alpha_{j}\right)=$ 4 .
@@
${\mathcal{L}}_{i}(\theta^{*})-\mathcal{L}_{i}(\theta_{i})\leq\frac{L_{i}C}{2}\left(1+2\alpha_{i}+\sum\limits_{j=1}^{T}\alpha_{j}^{2}\right)+\frac{L_{i}\epsilon C}{2}\left(\sum_{j\neq k}\alpha_{j}\alpha_{k}+\sum\limits_{j\neq i}^{T}\alpha_{j}\right)\leq 2L_{i}C(1+\epsilon).\qed$

### v2

@@ 6.1 Models and Datasets
This paper includes both computer vision and natural language processing experiments. In the vision experiments, we use CLIP (Radford et al., 2021) and ViT (Dosovitskiy, 2020) models from OpenCLIP (Ilharco et al., 2021), along with the following 8 datasets: MNIST (LeCun et al., 1998), EuroSAT (Helber et al., 2019), Cars (Krause et al., 2013), GTSRB (Stallkamp et al., 2012), RESISC45 (Cheng et al., 2017), DTD (Cimpoi et al., 2014), SUN397 (Xiao et al., 2010), and SVHN (Netzer et al., 2011). The language experiments are based on RoBERTa (Liu, 2019) models, tested on a medium scale benchmark with 12 tasks (Panigrahi et al., 2023; He et al., 2024) and a larger scale collection of 70 tasks, the latter with further details provided in Appendix D.
@@ 6.2 Empirical Verification of Theoretical Results
We primarily focus on 4.4 and first examine the relationship between the loss gap and key constants. Then, we report the empirical loss gap to support our upper bounds. All experiments in this section imply that our theoretical analysis in Section 4.2 provides actionable insights that inform the practical design of new merging algorithms: one should minimize the task vector norm $C$, similarity $\epsilon$, and smoothness $L$ for alignment to minimize the upper bound of the loss gap for better generalization performance.
6.2.1 Role of Constants in Loss Gap
@@
To verify how task vector similarity $\epsilon$ impacts the performance, we conduct the experiment shown in Figure 3. We merge the MNIST task vector with each of the other task vectors, all having similar norms ranging from $[2,3)$ (see Table 9 in the Appendix for details), and set the scaling coefficient $\alpha$ to $0.5$. In this setting, we approximately control all constants in 4.4, including $L$, $\alpha$, and $C$, and observe that highly similar tasks, such as digit classification in MNIST, SVHN, and GTSRB, lead to larger loss gaps or worse performance for MNIST compared to less related tasks.
@@
We provide additional evidence that both $C$ and $\epsilon$ must be constrained for the success of task vector addition. In Table 2, we collected task vectors for 8 tasks from two CLIP checkpoints pretrained with different hyperparameters. From this table, we observe that successful task addition reveals the identity of the task vectors. For optimal merging performance, we should only add task vectors fine-tuned from the same checkpoint, as any mixture of task vectors from different checkpoints will cause a significant performance drop. The above empirical observation consolidates our 4.1 that task vectors should reside in the same fine-tuning regime. To elaborate, from Figure 4, although all $\epsilon$ values in the pink box are very low, task addition still fails due to the large $C$ value. From Figure 5, we see that with different hyperparameters, the two pretrained models are situated in two local convex basins, and the distance between the two checkpoints is much larger than the task vectors (for more discussion, see Section C.1). Thus, if we create task vectors by subtracting the wrong pretrained checkpoint, the large $C$ value leads to the failure of task addition.
@@
6.2.2 Loss Gap in Task Arithmetics
We report the empirical loss gap using Task Arithmetic as the merging method on several downstream datasets using the ViT-B/16 backbone for CLIP. In Table 3, the loss gap is computed as the empirical difference between the loss of the merged model $\mathcal{L}_{i}(\theta^{*})$ and the fine-tuned loss $\mathcal{L}_{i}(\theta_{i})$ as in the left hand side of theorems in Section 4.2.
Table 3: Empirical loss gap between the fine-tuned model and the TA-merged model for each task on ViT-B/16.
Taski $\mathcal{L}_{i}(\theta^{*}_{\text{TA}})-\mathcal{L}_{i}(\theta_{i})$
MNIST 0.12
EuroSAT 1.00
Cars 0.15
DTD 1.11
GTSRB 0.80
RESISC45 0.84
SUN397 5.02
SVHN 0.60
As shown in Table 3, the loss gap remains small for most tasks except SUN397, which reaches a gap of $5.02$. This demonstrates that although task arithmetic (Ilharco et al., 2022) is a widely used model merging baseline, it can suffer significant performance degradation on certain tasks. Importantly, these empirical loss gaps are not computed from the constants appearing in the right-hand side of our theoretical upper bound. When substituting reasonable values, such as $L=1$, $C=4$, and $\epsilon=0.2$, the upper bound becomes $LC(1+\epsilon)\approx 5$, which approximates the largest observed gap. While this value may seem conservative for common losses like cross-entropy, it is necessary to account for the worst-case scenarios observed in practice. This further supports the relevance of our theoretical framework in explaining and bounding the limitations of task arithmetic in heterogeneous settings.
@@ 6.3 Vision Experiments with Task Vector Bases
Table 5: Accuracy of different merging strategies (Save One, MS-ER, Save All) using the same algorithm as $\texttt{merge}_{\text{in}}$ and $\texttt{merge}_{\text{out}}$ under continual setting, averaged across ViT architectures.
merge Save One MS-ER Save All
Mean 86.04 89.88 89.88
TIES 84.58 91.41 92.48
Topk 73.12 73.63 72.61
@@
For a detailed analysis of the behavior of task vector bases methods, we first fix a small subset of tasks where the separation between digit classification (MNIST, SVHN) and satellite image classification (EuroSAT, RESISC45) is obvious and matches the optimal clustering result that reduces the memory storage by half. With only $4$ task vectors, we can enumerate all possible 15 partitions of tasks and understand the position of the merging performance using this natural partition. Note that scaling this performance analysis to 8 vision tasks would require analyzing 4,140 partitions, which is computationally infeasible. We provide only clustering results for the 8-dataset benchmark in Section C.2, and we see MNIST–SVHN and EuroSAT–RESISC45 pairings remain consistent.
@@
After comparing the base addition performance with the average of all partitions in Table 4, next in Table 5, we repeat the addition experiments with the same inner and outer merging methods to validate the scenario illustrated in Figure 2. Assuming any of M, S, E, or R as the next incoming task vector, Save One corresponds to the average over four 3-1 partitions (ex., MSE-R), while Save All corresponds to $k=4$ (M-S-E-R) in Table 12 in the Appendix. As expected, saving a single model yields the lowest accuracy, while using bases achieves better performance than the single model and yields the highest accuracy among the three scenarios.
@@
After saving task vector bases from MNIST, SVHN, EuroSAT, and RESISC45 in Section 6.3, we test generalization metrics on unseen tasks in our vision benchmark. We set a similarity threshold to only use the most similar task vector bases for GTSRB and DTD generalization in Table 7, with $\texttt{merge}_{\text{out}}$ set to Task Arithmetic. Since we use only one basis for generalization, this reduces to tuning $\alpha$ for the selected basis on the validation set. For GTSRB, which involves traffic sign recognition, we expect overlap with the digit classification skills in MNIST and SVHN. In Table 7, we observe that the performance of the MS-ER column, corresponding to the bases method, is generally higher than the M-S-E-R column, where full task arithmetic flexibility is retained. For OOD generalization, using all task-specific information related to the target task proves beneficial. However, since DTD is not related to any stored bases, its OOD generalization performance is almost equivalent to the pretrained model. This validates 4.5 and shows that when an OOD task is not too far away from seen tasks, leveraging bases knowledge is more effective than relying solely on a task vector trained on a single task, with the added benefit of memory savings.
@@ 6.4.1 Results for Large Scale Benchmark
We first fix 70 natural language understanding tasks from top downloaded Huggingface datasets (Wolf, 2019) with the tag of text classification and multiple choice questions datasets that can be reformulated as classification problems. With a larger number of tasks, we can conduct a stress test to determine how much we can compress the memory budget using Algorithm 1. We reuse prompt templates and labels in Gao et al. (2020); He et al. (2024) for masked language modeling fine-tuning of RoBERTa-base, ensuring that each fine-tuned model’s performance exceeds the majority vote baseline. As a result, each of the collected models is meaningful, and since all tasks can be evaluated with accuracy, the merged metrics are directly comparable.
We first examine the result of clustering in Figure 7, where we show the task cosine similarity matrix ordered by the spectral clustering results. We plot two matrices: $k=3$, which represents the number of clusters selected based on the largest eigenvalue gap, and $k=10$, which further partitions tasks into more fine-grained groups. Consequently, The properties of the task vectors change according to the prompt template difference. When $k=3$, the tasks are split based on the number of input sentences, and when $k=10$, there are more fine-grained clusters that reflect the task distribution, such as the separation of bAbI-NLI tasks, which reformulate tasks from Weston et al. (2015) as NLI problems, and other types of NLI tasks. In summary, the generated task clusters in step 2 of Figure 1 are reasonable, as they reflect certain interpretable skills learned by all models. And when we use $k=3$ as the algorithm suggested, we reduce the storage of 70 task vectors into 3 bases, saving 95.71% of the memory budget. In Section D.4, we further show that such grouping is the result of inherent task properties, not just due to prompt design.
Table 8: Merged accuracy/F1 and storage cost on the $12$-task language benchmark. First two rows contain numbers in He et al. (2024). For each method, we compute: (1) #Masks used (for non-sparsity based methods, this is equal to the number of tasks in the benchmark), and (2) Sparsity%, which is the percentage of non-zero entries averaged across masks. The total storage cost is then computed as: #Masks $\times$ %Sparsity, which can be seen as the size of memory bank in the setup of Figure 2. Our Mask-Sharing version of L&S achieves the best overall memory metric. Stitch-Twice version of L&S improves upon all previous baselines (non-shaded rows, and more competitive baselines see Section D.1) in both accuracy and memory usage.
@@ 6.4.2 Results for Medium Scale Benchmark
(L&S) He et al. (2024) divides model merging into two steps: first, localizing the task information by learning a sparse vector mask, and second, stitching sparse models with mask normalization, ensuring that $\alpha$ is a probability vector like 4.3. Compared to other model merging methods, He et al. (2024) offers the dual advantage with significantly improved multitask performance due to reduced conflict from localization, and drastically lowered memory costs by saving only a small proportion of nonzero weights. Thus, to maximally boost all metrics, we investigate how to integrate L&S into our bases creation framework which requires two steps of merging. As described in Section 5.1, after clustering sparse localized task vectors, we can directly apply L&S as $\texttt{merge}_{\text{in}}$ within each cluster. For $\texttt{merge}_{\text{out}}$, it requires special design and we propose two modifications to L&S.
@@
The second approach, Stitch Twice, first localizes tasks and stitches by mask normalization within cluster. Afterward, in $\texttt{merge}_{\text{out}}$, we round within-cluster-normalized masks as updated binary task masks, and use them as input for the second iteration of stitching across all $T$ tasks. Formally, in the first iteration, for $l\in[d]$, the $l$-th entry of the mask for task $i$ is $\gamma_{i}^{l}\leftarrow\texttt{round}(\gamma_{i}^{l}/\sum_{j=1}^{N}\gamma_{j}^{l})$, and the second iteration of mask update is $\gamma_{i}^{l}\leftarrow\gamma_{i}^{l}/\sum_{j=1}^{T}\gamma_{j}^{l}$. Inspired by the improvements shown in Figure 8, this two-step process automatically readjusts task vector weights with appropriate task relationships learned from clustering. Unlike Mask Sharing, Stitch Twice requires $T$ copies of sparse models, since each task retains its own mask. However, in Table 8, Stitch Twice improves generalization performance from L&S while also surprisingly reducing memory costs due to increased mask sparsity from double normalization.
@@ 7 Conclusion and Future Work
In this work, we revisited task vectors as a practical approach for model editing. By providing a theoretical foundation for task arithmetic, we addressed key gaps in understanding its underlying mechanisms and identified the conditions for its success. Building on theoretical insights, we introduced a scalable framework that reduces the memory footprint of task vector arithmetic through clustered task bases, significantly improving efficiency while maintaining high performance, and demonstrated the flexibility of bases arithmetic. Looking forward, we foresee several future directions. First, the creation of our task vector bases depends on the task relationship, and there are additional opportunities to improve memory efficiency from hardware perspectives. Additionally, by clustering, we assume that there are disjoint groups of tasks for the initial task vectors. Therefore, it might be interesting to further decompose and reorganize task vectors to boost performance for more complicated task relationships.
@@ Acknowledgment
Part of this work was done when both SZ and HZ were visiting the Simons Institute via the Modern Paradigm in Generalization program. SZ, YH, and HZ are partially supported by an NSF IIS grant No. 2416897. HZ would like to thank the support from a Google Research Scholar Award. WY is supported by a gift from AWS AI for Trustworthy AI, as well as funding from the Defense Advanced Research Projects Agency’s (DARPA) SciFy program (Agreement No. HR00112520300). MY is partly supported by MEXT KAKENHI Grant Number 24K03004 and by JST ASPIRE JPMJAP2302. We want to express a special thanks to Yuechun Sun who pointed out a way to improve the upper bound for task addition theorem. We are also thankful to Wei Hu, Yuki Takezawa, Jose Renato Restom and anonymous reviewers for other helpful discussions. The views and conclusions expressed in this paper are solely those of the authors and do not necessarily reflect the official policies or positions of the supporting companies and government agencies.
@@
which means that $\theta^{*}$ is within the fine-tuning regime and satisfies the local smoothness assumption. Hence, if $x\sim\mathcal{D}_{i}$
${\mathcal{L}}_{i}(\theta^{*})-\mathcal{L}_{i}(\theta_{i})$ $\theta^{*}-\theta_{i},\frac{\partial\mathcal{L}_{i}(x,\theta_{i})}{\partial{\theta}}\right>+\frac{L_{i}}{2}\left\lVert\theta^{*}-\theta_{i}\right\rVert^{2}$ (4.2)
@@
To bound the second norm term, we reassign the subscript $\alpha_{k}:=\alpha_{i}$ as the coefficient for the $i$-th task to avoid confusion with the summation indices. Next we define $t_{j}:=\alpha_{j}\ (j\neq k),t_{k}:=\alpha_{k}-1$. Since $\sum_{i}\alpha_{i}=1$, we have $t_{k}+\sum_{j\neq k}t_{j}=0$. Then,
$\left\lVert\sum_{j=1}^{T}\alpha_{j}\tau_{j}-\tau_{i}\right\rVert^{2}$ $=\left\lVert\sum_{j=1}^{T}\alpha_{j}\tau_{j}-\tau_{k}\right\rVert^{2}$
$=\left\lVert\sum_{i=1}^{n}t_{i}\tau_{i}\right\rVert^{2}$
$\leq\sum_{i=1}^{n}{t_{i}}^{2}\tau_{i}^{2}+2\sum_{1\leq i<j\leq n}\left|t_{i}t_{j}\right|\cdot\left|\left<\tau_{i},\tau_{j}\right>\right|$
$\leq C\left(\sum_{i=1}^{n}t_{i}^{2}+2\epsilon\sum_{1\leq i<j\leq n}\left|t_{i}t_{j}\right|\right)$ (1)
By 4.3, $t_{k}\leq 0$. Besides, we have $-t_{k}=t_{1}+\dots+t_{k-1}+t_{k+1}+\dots+t_{n}=\sum_{i\neq k}t_{i}$. We can bound the cross-product term as follows:
$\sum_{1\leq i<j\leq n}\left|t_{i}t_{j}\right|$ $=-t_{k}\sum_{i\neq k}\left|t_{i}\right|+\sum_{1\leq i<j\leq n,i\neq k,j\neq k}t_{i}t_{j}$
$=t_{k}^{2}+\sum_{1\leq i<j\leq n,i\neq k,j\neq k}t_{i}t_{j}$
$=t_{k}^{2}+\frac{1}{2}\left(\left(\sum_{i\neq k}t_{i}\right)^{2}-\sum_{i\neq k}t_{i}^{2}\right)$
$=t_{k}^{2}+\frac{1}{2}\left(t_{k}^{2}-\left(\sum_{i}t_{i}^{2}-t_{k}^{2}\right)\right)$
$=2t_{k}^{2}-\frac{1}{2}\sum_{i}t_{i}^{2}$
Plug it back into Equation 1, we have
$\sum_{i}t_{i}^{2}+2\epsilon\sum_{1\leq i<j\leq n}\left|t_{i}t_{j}\right|$ $=\sum_{i}t_{i}^{2}+2\epsilon\left(2t_{k}^{2}-\frac{1}{2}\sum_{i}t_{i}^{2}\right)$
$=(1-\epsilon)\sum_{i}t_{i}^{2}+4\epsilon t_{k}^{2}$
$=(1-\epsilon)\left[\sum_{i\neq k}\alpha_{i}^{2}+(\alpha_{k}-1)^{2}\right]+4\epsilon(\alpha_{k}-1)^{2}$
$=(1-\epsilon)\left[\sum_{i}\alpha_{i}^{2}-2\alpha_{k}+1\right]+4\epsilon(\alpha_{k}-1)^{2}$
$\leq(1-\epsilon)\cdot 2+4\epsilon\cdot 1$ (4.3)
$=2+2\epsilon$
@@
${\mathcal{L}}_{i}(\theta^{*})-\mathcal{L}_{i}(\theta_{i})\leq\frac{L_{i}C}{2}(2+2\epsilon)=L_{i}C(1+\epsilon).\qed$